OpenStudy (anonymous):

Solve for the variable in the equation 14 = t + 12

OpenStudy (anonymous):

subtract 12 from both sides

OpenStudy (anonymous):

\[14=t+12\] \[14-12=t+12-12\] \[2=t\]

OpenStudy (anonymous):

probably just easier to think "what would i add to 12 to get 14?" answer: 2

OpenStudy (anonymous):

O i get it now

OpenStudy (anonymous):

good !